OpenStudy (anonymous):

I need help with transformations of functions. There's two problems. Describe the transformations of the function f(x)=3 divided by x+2, +4 Describe the transformations of the function f(x)=-2 divided by x-1

OpenStudy (danjs):

for the first.. f(x) = y = 3 that just describes a horizontal line right?

OpenStudy (danjs):

oh i see what the question is asking now.. \[y = \frac{ 3 }{ x+2 }\]

OpenStudy (danjs):

looks like it wanted you to compare y=3 to y=3/(x+2) from the wording

OpenStudy (danjs):

In general if you replace x, by (x - a), the graph will be shifted horizontally by a

OpenStudy (danjs):

For example, Transforming y = 3/x to y=3/(x + 3) will shift the graph a = -3 units

OpenStudy (jdoe0001):

hmm can you post a quick screenshot of the material? so we can see what they mean

OpenStudy (jdoe0001):

looks rather ambiguous, at least to me

OpenStudy (danjs):

I think it is just describing the difference in \[f(x) = \frac{ 3 }{ x }~~~~and~~~~f(x) = \frac{ 3 }{ x+3 }\]

OpenStudy (danjs):

then 3/(x+4)

OpenStudy (danjs):

If you replace x by (x - a) the graph shifts horizontally by 'a' units. \[f(x) = \frac{ 3 }{ x-a }\] in the first, a = -3, the graph moves 3 units in the negative direction, the others are the same, just with different numbers.,

OpenStudy (danjs):

remember (x-a) = (x + (-a))